# Robust Stochastically-Descending Unrolled Networks

### v1

@@ V-A Problem Formulation
$\begin{split}\arg\min_{{\bf z}}\quad&\|{\bf x}-{\bf A}g_{\mu}({\bf z})\|_{2}^{2}+\lambda\|{\bf z}\|_{2}^{2},\end{split}$ (19)
where $\lambda$ is a regularization parameter balancing the data consistency and the prior. Finding the optimal of (19) requires the generative model $g_{\mu}$ to be known apriori. This implies that learning the parameters of $g_{\mu}$ is executed separately before (19) is solved.
@@
Figure 6: Comparison between constrained and unconstrained GLOW-Prox. (Left) Distance to the original image $\bf y$, averaged over the test dataset. (Right) The $\ell_{2}$-norm of the latent variable ${\bf z}_{l}$, representing the negative log-likelihood.
@@ V-B Unrolled GLOW-Prox
where $\zeta_{l}$ is a learnable shrinkage parameter. The proximal update ensures that the estimate has a high likelihood by getting ${\bf z}_{l}$ closer to the origin, i.e., reducing its $\ell_{2}$-norm. Finally, we convert back to the signal space using the inverse map $g_{\mu}^{(l)}$,
@@ V-C Constrained GLOW-Prox
Unlike [7], we perform a constrained-learning training procedure according to (CO). The unrolled network is trained by minimizing the mean square error between the output of unrolled GLOW-Prox and the true image,
[…]
Here, we use two sets of the gradient constraints (CI) as we separate the variables ${\bf y}$ and ${\bf z}$ when computing the gradients, following the separation of their update rules in (20) and (22). In the first set of constraints, we force the gradient of the data consistency term with respect to ${\bf y}$, or $g_{\mu}({\bf z})$, to descend, thereby ensuring that the update rule of ${\bf z}$ in (22) does not drift the estimate ${\bf y}_{l}$ away from $\tilde{\bf y}_{l}$. In other words, $\tilde{\bf y}_{l}$ is, by definition, chosen in the direction of the gradient while ${\bf y}_{l}$ is forced to be in a descending direction using the imposed constraints. In the second set, we consider the gradients of the regularization term with respect to the latent variable ${\bf z}$ to ensure that its likelihood is increasing over the layers. The training procedure follows Algorithm 1 except that two sets of dual variables are considered to accompany the two sets of constraints.
Figure 7: Test examples of inpainted images across the unrolled layers of (left) constrained and (right) unconstrained GLOW-Prox. The performance of the two appears to be almost identical.
@@ V-D Numerical Results and Discussions
The unrolled GLOW-Prox consists of $L=5$ layers. Each layer contains a GLOW network that has a depth of flow of $18$ and $4$ multi-scale levels [47]. We refer the reader to [7, Fig. 1] for a visualization of the network. We train the unrolled model to minimize (24) using ADAM with a learning rate $\mu_{w}=10^{-5}$ and a dual learning rate $\mu_{\lambda}=10^{-4}$. The constraint parameter $\epsilon$ is set to $0.05$. As per (CO), we add a noise vector ${\bf n}_{l}\sim{\cal N}(0,\sigma^{2}_{l}{\bf I})$ to the output of each unrolled layer $l$ to ensure that the constraints are statistically independent. The noise variance declines over the layers, that is, $\sigma^{2}_{l}\propto\frac{1}{l}$. The training is executed with mini-batches of size $8$ and was held for $30$ epochs. [^2] The code is available at: https://github.com/SMRhadou/UnrolledGlow. We compare the performance and robustness of our constrained GLOW-Prox to the standard version trained without constraints [7].
@@
Performance. We evaluate the distance, averaged over the test dataset, between the output of each unrolled layer and the clean images. The results are reported in Figure 6 (left), which shows that both constrained and unconstrained GLOW-Prox have a similar behavior in the image domain. This is also confirmed in Figure 7, which displays three test examples. As shown in the figure, the networks gradually inpaint the masked area reflecting a descending nature in the layers’ estimates. The two methods, however, depart from each other in the latent space, as shown in Figure 6 (right). The $\ell_{2}$-norm of the latent variable decreases gradually when the network is trained using constraints, implying that the gradient constraints were satisfied (and generalized) in the test examples. This is not the case when we omit the constraints during training, as depicted by the red line that fluctuates over the layers. However, we observe a gap in the $\ell_{2}$-norm between the two models, with the constrained one achieving higher norms across all the layers. This can be attributed to the noise added to the layers in the constrained case, which adds uncertainty to the network and reduces the log-likelihood.
Robustness. Despite the aforementioned slight drawback, constrained GLOW-Prox shows more robustness to input perturbations, as illustrated in Figure 8. The left plot in the figure shows the MSE between the original and reconstructed images after adding perturbations to the input image. The perturbation is a random matrix $\frac{p{\bf N}}{\|{\bf N}\|_{2}}$, where $\bf N$ is sampled from a standard normal distribution and $p$ represents the perturbation size. As can be inferred from Figure 8 (left), the performance of the two models deteriorates with the perturbation size; however, GLOW-Prox deteriorates faster and more significantly than the constrained model. The distinction between the two models is more perceptible in Figure 8 (right), where the distortion in the shape of the facial features is more pronounced in the case of standard GLOW-Prox. In the case of constrained GLOW-Prox, the reconstructed images, albeit slightly different from the original, are still reasonably acceptable even under severe perturbations sizes.
@@
Figure 8: Robustness. (Left) Mean square error between the original and reconstructed (inpainted) images under different perturbation sizes $p$. (Right) Examples of inpainted images using (top) constrained GLOW-Prox and (bottom) unconstrained GLOW-Prox for different $p$. Deterioration in the quality of the reconstructed images occurs faster in the unconstrained case.
@@ VI Conclusions
In this paper, we introduced a framework for posing unrolled networks as stochastic descent algorithms. Within this framework, the unrolled layers are trained with descending constraints to ensure that their trajectory of estimates descends toward a stationary point of the optimizee problem. Our theoretical analysis shows that this trajectory is guaranteed to converge with an exponential rate of convergence. These substantiated convergence guarantees confer two pivotal advantages upon unrolled networks, as demonstrated through our simulations. Firstly, the unrolled networks are guaranteed to generalize to previously unseen in-distribution problems. Secondly, they exhibit robustness to additive perturbations, thereby enhancing their out-of-distribution generalizability. These findings collectively underscore the potential of the proposed framework to contribute to the development of robust, interpretable unrolled networks for a wider array of applications.
@@ Appendix A Proof of Theorem 2
We then define a sequence of random variables with a degenerate distribution such that ${Z_{l}}={\mathbb{E}\|{\nabla}f({\bf y}_{l})\|}$ almost surely for all $l$. We then construct a supermartingale-like inequality from (26):
$\begin{split}&\mathbb{E}[Z_{l}|\ {\cal F}_{l-1}]\leq(1-\delta)(1-{\epsilon})\ Z_{l-1}+(1-\delta)\zeta(N,\delta)+\delta M\\ &=(1-\delta)\ Z_{l-1}-(1-\delta)\Big{(}\epsilon Z_{l-1}-\zeta(N,\delta)-\frac{\delta M}{1-\delta}\Big{)}.\end{split}$ (27)
The goal of the rest of the proof is to show that i) when $l$ grows, $Z_{l}$ almost surely and infinitely often achieves values below $\eta:=\frac{1}{\epsilon}\big{(}\zeta(N,\delta)+\delta M/1-\delta\big{)}$, and that ii) this also holds for the gradient norm $\|{\nabla}f({\bf y}_{l})\|$ itself.
@@
It is worth noting that, for all $l$, $\alpha_{l}\geq 0$ and $\beta_{l}\geq 0$ by construction. It then follows from supermartingale convergence theorem [51, Theorem 1] that (33) implies that (i) $\alpha_{l}$ converges almost surely, and (ii) $\sum_{l=1}^{\infty}\beta_{l}$ is almost surely summable (i.e., finite). When the latter is written explicitly, we get
@@
By Fatou’s lemma [52, Theorem 1.5.5], it follows that

### v2

@@ V-A Problem Formulation
$\begin{split}\arg\min_{{\bf z}}\quad&f_{in}({\bf z};{\bf x}):=\|{\bf x}-{\bf A}g_{\mu}({\bf z})\|_{2}^{2}+\lambda\|{\bf z}\|_{2}^{2},\end{split}$ (21)
where $\lambda$ is a regularization parameter balancing the data consistency and the prior. Finding the optimal of (21) requires the generative model $g_{\mu}$ to be known apriori. This implies that learning the parameters of $g_{\mu}$ is executed separately before (21) is solved. We denote the data consistency term $\|{\bf x}-{\bf A}{\bf y}\|_{2}^{2}$ by $f_{1}({\bf y})$ and the regularization term $\|{\bf z}\|_{2}^{2}$ by $f_{2}({\bf z})$.
@@ V-B Unrolled GLOW-Prox
where $\zeta_{l}$ is a learnable shrinkage parameter. The proximal update ensures that the estimate has a higher likelihood–equivalent to a lower $\ell_{2}$ norm since $\bf z$ follows a standard normal distribution–by bringing ${\bf z}_{l}$ closer to the origin. Finally, we convert back to the signal space using the inverse map $g_{\mu}^{(l)}$,
@@ V-C Constrained GLOW-Prox
Unlike [7], we perform a constrained training procedure according to (CO). The unrolled network is trained by minimizing the mean square error between the output of unrolled GLOW-Prox and the true image, i.e.,
$\begin{split}\min_{{\bf W}}\quad&\mathbb{E}\big{[}\|\bm{\Phi}({\bf x};{\bf W})-{\bf y}\|_{2}^{2}\big{]}\\ \text{s.t.}\quad&\mathbb{E}\big{[}\|{\nabla}_{y}f_{1}({\bf y}_{l})\|_{2}-(1-\epsilon)\ \|{\nabla}_{y}f_{1}({\bf y}_{l-1})\|_{2}\big{]}\leq 0\\ &\mathbb{E}\big{[}\|{\nabla}_{z}f_{2}({\bf z}_{l})\|_{2}-(1-\epsilon)\ \|{\nabla}_{z}f_{2}({\bf z}_{l-1})\|_{2}\big{]}\leq 0,\end{split}$ (26)
where $f_{1}$ and $f_{2}$ are the data consistency and the regularization terms specified in (21). Here, we use two sets of the gradient constraints (CI) as we separate the variables ${\bf y}$ and ${\bf z}$ when computing the gradients, following the separation of their update rules in (22) and (24). In the first set of constraints, we force the gradient of the data consistency term $f_{1}({\bf y})$ with respect to ${\bf y}$ to descend, thereby ensuring that the update rule of ${\bf z}$ in (24) does not drift the estimate ${\bf y}_{l}$ away from $\tilde{\bf y}_{l}$. In other words, $\tilde{\bf y}_{l}$ is, by definition, chosen in the direction of the gradient while ${\bf y}_{l}$ is forced to be in a descending direction using the imposed constraints. In the second set, we consider the gradients of the regularization term $f_{2}({\bf z})$ with respect to the latent variable ${\bf z}$ to ensure that its likelihood is increasing over the layers. The training procedure follows Algorithm 1 except that two sets of dual variables are considered to accompany the two sets of constraints.
Figure 7: Comparison between constrained and unconstrained GLOW-Prox. (Left) Distance to the original image $\bf y$, averaged over the test dataset. (Middle) The $\ell_{2}$-norm of the latent variable ${\bf z}_{l}$, representing the negative log-likelihood. (Right) The value of the inpainting objective function $f_{in}(\cdot;{\bf x})$.
Figure 8: Test examples of inpainted images ${\bf y}_{l}$ across the unrolled layers of (left) constrained and (right) unconstrained GLOW-Prox. The performance of the two appears to be almost identical. The initial estimate ${\bf y}_{0}$ is set to be the masked image $\bf x$.
@@ V-D Numerical Results and Discussions
The unrolled GLOW-Prox consists of $L=5$ layers. Each layer contains a GLOW network that has a depth of flow of $18$ and $4$ multi-scale levels [50]. We refer the reader to [7, Fig. 1] for a visualization of the network. We train the unrolled model to minimize (26) using ADAM with a learning rate $\mu_{w}=10^{-5}$ and a dual learning rate $\mu_{\lambda}=10^{-3}$. The constraint parameter $\epsilon$ is set to $0.05$. As per (CO), we add a noise vector ${\bf n}_{l}\sim{\cal N}(0,\sigma^{2}_{l}{\bf I})$ to the output of each unrolled layer $l$ and the initial estimate ${\bf y}_{0}$ is set to be the masked image $\bf x$. The noise variance declines over the layers, that is, $\sigma^{2}_{l}\propto\frac{1}{l}$. The training is executed with mini-batches of size $8$ and was held for $30$ epochs. [^2] The code is available at: https://github.com/SMRhadou/UnrolledGlow. We compare the performance and robustness of our constrained GLOW-Prox to the standard version trained without constraints [7].
@@
Performance. We evaluate the distance, averaged over the test dataset, between the output of each unrolled layer and the clean images. The results are reported in Figure 7 (left), which shows that both constrained and unconstrained GLOW-Prox have a similar behavior in the image domain. This is also confirmed in Figure 8, which displays three test examples. As shown in the figure, the networks gradually inpaint the masked area reflecting a descending nature in the layers’ estimates. The two methods, however, depart from each other in the latent space, as shown in Figure 7 (middle). The $\ell_{2}$-norm of the latent variable decreases gradually when the network is trained using constraints, implying that the gradient constraints were satisfied (and generalized) in the test examples. This is not the case when we omit the constraints during training, as depicted by the red line that fluctuates over the layers. Combined together, Figure 7 (right) illustrates that the objective function $f_{in}$ indeed decreases over the layers in the case of constrained GLOW-prox.
@@
Figure 9: OOD Robustness. (Left) Mean square error between the original and reconstructed (inpainted) images under different perturbation sizes $p$. (Right) Examples of inpainted images using (top) constrained GLOW-Prox and (bottom) unconstrained GLOW-Prox for different $p$. Deterioration in the quality of the reconstructed images occurs faster in the unconstrained case.
Robustness. Constrained GLOW-Prox also shows more OOD robustness to input perturbations, as illustrated in Figure 9. The left plot in the figure shows the MSE between the original and reconstructed images after adding perturbations to the input image. The perturbation is a random matrix ${\bf N}$, where each element in the matrix is sampled iid sampled from a normal distribution ${\cal N}(0,p^{2})$ and $p$ represents the perturbation size. As can be inferred from Figure 9 (left), the performance of the two models deteriorates with the perturbation size; however, GLOW-Prox deteriorates faster and more significantly than the constrained model. The distinction between the two models is more perceptible in Figure 9 (right), where the distortion in the shape of the facial features is more pronounced in the case of standard GLOW-Prox. In the case of constrained GLOW-Prox, the reconstructed images, albeit slightly different from the original, are still reasonably acceptable even under severe perturbations sizes.
@@ VI Conclusions
In this paper, we introduced a framework for posing unrolled networks as stochastic descent algorithms. Within this framework, the unrolled layers are trained with descending constraints to ensure that their trajectory of estimates descends toward a stationary point of the optimizee problem. Our theoretical analysis shows that this trajectory is guaranteed to converge with an exponential rate of convergence. These substantiated convergence guarantees confer an advantage upon unrolled networks as we analytically show that the learnt descending behavior is transferable under distribution shifts in the optimization problem being solved. These findings collectively underscore the potential of the proposed framework to contribute to the development of robust, interpretable unrolled networks for a wider array of applications. One potential area for future work is the expansion of our framework to enforce the feasibility of the solutions in the context of constrained optimizee problems, which encompass many engineering problems.
@@ Appendix A Proof of Theorem 2
We let ${Z_{l}}={\mathbb{E}_{D_{x}}\|{\nabla}f({\bf y}_{l})\|}$ be the gradient norm averaged over the input data distribution $D_{x}$, which is a random variable depending on the value of ${\bf y}_{l-1}$ and ${\bf n}_{l}$. We then construct a supermartingale inequality from (28):
$\begin{split}&\mathbb{E}_{{\cal N}_{l}}[Z_{l}|\ {\cal F}_{l-1}]\leq(1-\delta)(1-{\epsilon})\ Z_{l-1}+(1-\delta)\zeta(N,\delta)+\delta M\\ &=(1-\delta)\ Z_{l-1}-(1-\delta)\Big{(}\epsilon Z_{l-1}-\zeta(N,\delta)-\frac{\delta M}{1-\delta}\Big{)}.\end{split}$ (29)
The conditional expectation on the left-hand side is taken over the distribution of the noise ${\bf n}_{l}$. The goal of the rest of the proof is to show that i) when $l$ grows, $Z_{l}$ almost surely and infinitely often achieves values below $\eta:=\frac{1}{\epsilon}\big{(}\zeta(N,\delta)+\delta M/1-\delta\big{)}$, and that ii) this also holds for the gradient norm $\|{\nabla}f({\bf y}_{l})\|$ itself.
@@
It is worth noting that, for all $l$, $\alpha_{l}\geq 0$ and $\beta_{l}\geq 0$ by construction. It then follows from supermartingale convergence theorem [54, Theorem 1] that (35) implies that (i) $\alpha_{l}$ converges almost surely, and (ii) $\sum_{l=1}^{\infty}\beta_{l}$ is almost surely summable (i.e., finite). When the latter is written explicitly, we get
@@
By Fatou’s lemma [55, Theorem 1.5.5], it follows that
@@ Appendix B Proof of Lemma 1
Appendix C Proof of Corollary 1
Proof.
We start by adding and subtracting $\mathbb{E}_{D_{x}}\big{[}\|{\nabla}f({\bf y}_{l};{\bf x})\|_{2}\big{]}$ and $(1-\epsilon)\Big{[}\mathbb{E}_{D_{x}}\big{[}\|{\nabla}f({\bf y}_{l-1};{\bf x})\|_{2}]$ from the quantity we seek to evaluate, i.e., we get
$\begin{split}&\ \mathbb{E}_{D_{x}^{\prime}}\big{[}\|{\nabla}f({\bf y}_{l};{\bf x})\|_{2}]-(1-\epsilon)\ \mathbb{E}_{D_{x}^{\prime}}\big{[}\|{\nabla}f({\bf y}_{l-1};{\bf x})\|_{2}\big{]}\\ &=\mathbb{E}_{D_{x}^{\prime}}\big{[}\|{\nabla}f({\bf y}_{l};{\bf x})\|_{2}]-\ \mathbb{E}_{D_{x}}\big{[}\|{\nabla}f({\bf y}_{l};{\bf x})\|_{2}\big{]}\\ &+(1-\epsilon)\Big{[}\mathbb{E}_{D_{x}}\big{[}\|{\nabla}f({\bf y}_{l-1};{\bf x})\|_{2}]-\ \mathbb{E}_{D_{x}^{\prime}}\big{[}\|{\nabla}f({\bf y}_{l-1};{\bf x})\|_{2}\big{]}\Big{]}\\ &+\mathbb{E}_{D_{x}}\big{[}\|{\nabla}f({\bf y}_{l};{\bf x})\|_{2}]-(1-\epsilon)\ \mathbb{E}_{D_{x}}\big{[}\|{\nabla}f({\bf y}_{l-1};{\bf x})\|_{2}\big{]}.\end{split}$ (51)
The right-hand side consists of three terms that can be bounded above with positive quantities according to Assumption 6 and (8). Therefore, the
$\begin{split}\mathbb{E}_{D_{x}^{\prime}}\big{[}&\|{\nabla}f({\bf y}_{l};{\bf x})\|_{2}]-(1-\epsilon)\ \mathbb{E}_{D_{x}^{\prime}}\big{[}\|{\nabla}f({\bf y}_{l-1};{\bf x})\|_{2}\big{]}\\ &\leq Md(D_{x},D_{x}^{\prime})+M(1-\epsilon)d(D_{x},D_{x}^{\prime})+\zeta(N,\delta)\\ &\leq 2Md(D_{x},D_{x}^{\prime})+\zeta(N,\delta).\end{split}$ (52)
Notice that this inequality holds with probability $1-\delta$ since the upper bound in (8) also holds with the same probability. This completes the proof.